OpenStudy (anonymous):

f(x,y)= x^2-y^2+xy-10x+3 a.) critical point b.) maximun, minimum or saddle point

OpenStudy (anonymous):

fx= 2x+y-10 fxx= 2 fxy=1 fy=-2y+x fyy=-2 fy= -2y+x=0 ---> x=2y fx= 2x+y-10 --> 2(2y)+y-10=0 ---> 4y+y=10 ---> y=2 x=2y--> x= 2*2= 4 the point is (2,2) then search D D= fxx*fyy-(fxy)^2 = -4 -1 = -5, D<0, means saddle

OpenStudy (anonymous):

fx= 2x+y-10 fxx= 2 fxy=1 fy=-2y+x fyy=-2 fy= -2y+x=0 ---> x=2y fx= 2x+y-10 --> 2(2y)+y-10=0 ---> 4y+y=10 ---> y=2 x=2y--> x= 2*2= 4 the point is (2,2) then search D D= fxx*fyy-(fxy)^2 = -4 -1 = -5, D<0, means saddle

OpenStudy (anonymous):

u can find the other point